OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days. f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 2 to n = 6 and what does it represent? (4 points)

OpenStudy (anonymous):

I think you mean f(n) = 8(1.05)^n, a starting height of 8 and a growth of 5% per day for n days. 11.26 = 8 (1.05)^n will tell you how many days he studied it. log(11.26/8) = n log(1.05) solve to find n Perhaps, though this is just linear equation y = 8(1.05)(n) which would have a slope of 8.40 and intercept =0.

OpenStudy (anonymous):

what part s this ? @douglaswinslowcooper